(5) Policyholders of an insurance company are classified as either high-risk or low-risk, 29% of policyholders (from both groups collectively) did not file a claim. 20% of high-risk policy holders did not file a claim, and 50% of low-risk policyholders did not file a claim. Given that a policyholder did file a claim, what was the probability that he was classified as high-risk? A) 0.54 B) 0.61 0.65 D) 0.79 E) 0.83 The circled answer is wrong, please show a good work to arrive at a correct answer please.
Answer #1

From the given data,the following Table is formed:

File Did not file Total
High risk 0.71 - 0.15 = 0.56 0.20x=0.14 x=0.7
Low risk 0.3 - 0.15 = 0.15 0.29- 0.20x=0.50(1-x)=0.15 1 - x=0.3
Total 1 - 0.29 = 0.71 0.29 1.00

Step 1:

Let x be the total percentage of High risk.

To find x:

Given:

Total Did not file = 0.29

So,

Total File = 1 - 0.29 = 0.71

Total percentage of low risk = 1 - x

Given:

20% of High risk did not file.

So,

Entry in High risk & Did not file = 0.20x

Given:

50% of low risk did not file.

So

Entry in Low risk & Did not file = 0.50(1-x)

0.29 - 0.20 x = 0.50(1-x)

So,

0.30 x = 0.50 - 0.29

So,

x = 0.7

Step 2:

Since x = 0.7 is got all the entries in the Table are filled.

So,

P(High risk/ File) = P(High risk & File)/ P(File)

= 0.56/ 0.71 = 0.79

So,

Correct option:

D 0.79